OpenStudy (anonymous):

Would someone help me review my integration homework? Area between graphs is its focus.

OpenStudy (abdullah1995):

i can try.

OpenStudy (irishboy123):

looks good

OpenStudy (zehanz):

I guess you don't need any help :). All your answers are correct. I understand every step you wrote down. One (tiny) point of criticism: in problem 3, the answer can be simplified just a little bit more to \(\sqrt 2 -1\).

OpenStudy (anonymous):

@ZeHanz wolfram says that number 5 is incorrect, and it says it's 44/15, not 74/15, but I can't find the error

OpenStudy (amistre64):

what are your permissible means of finding th eintersection points?

OpenStudy (anonymous):

seemingly any, including process of elimination

OpenStudy (zehanz):

It is just an silly error in one of the last steps: -96-40=-136, so the answer becomes (180-136)/15=44/15..

OpenStudy (anonymous):

doh, thanks for checking, good to go, and thanks to all for help

OpenStudy (zehanz):

YW!

OpenStudy (amistre64):

just to work some algebra .... x^4 -2x^3 +x^2 -x = x x=0 is obvious, divide by x x^3 -2x^2 +x -1 = 1 x^3 -2x^2 +x -2 = 0 factor x^2(x -2) +x -2 = 0 factor again (x^2+x) (x -2) = 0 since x not = 0 so we could divide ... x=2 is valid